OpenStudy (anonymous):

What is the equation of the line that passes through ( 1,5) and (4,11)?

OpenStudy (anonymous):

options??

OpenStudy (anonymous):

a. y = 6 b. y = x + 2 c. y = 2x + 3 d. y = 3x + 4

OpenStudy (anonymous):

ok its D

OpenStudy (anonymous):

thank you again

OpenStudy (anonymous):

welcome